Gremlin Industries will pay a dividend of $ 1.35 per share this year. It is expected that this dividend will grow by 5% per year each year in the future. The current price of​ Gremlin's stock is $ 23.20 per share. What is​ Gremlin's equity cost of​ capital?

Can someone explain this please
Answer #1

Price = D1/(ke-g) = $23.20 D1 = $1.35 5.00% ke-g = 5.82% (1.35/23.2] Cost of equity = ke 10.82% [5% + 5.82% ]